OpenStudy (mouth4war):

Line A and line B are perpendicular. Line A contains the points (4,2) and (-1, 3). If one of the points on line B is (3, 7), find another point on line B.

OpenStudy (solomonzelman):

you need to find the equation of line B, and then choose any other point on it. The question we need to adress though, is how to find the equation of this line B?

OpenStudy (solomonzelman):

Can you find the slope of line A for me please?

OpenStudy (solomonzelman):

\(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2}}{\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2}} }\) where \(\LARGE \color{black}{ \displaystyle {\rm m} }\) is the slope \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points.

OpenStudy (mouth4war):

-0.2

OpenStudy (solomonzelman):

m = (3-2)/(-1-4) = 1/-5 = -0.2 Correct

OpenStudy (solomonzelman):

Lets use a fraction -1/5 for now

OpenStudy (solomonzelman):

And you know that when a line C and D are prependicular to each other and line C has a slope \(\large\color{black}{ \displaystyle \rm M }\) , then the slope of line D must be \(\large\color{black}{ \displaystyle \rm -1/M }\).

OpenStudy (solomonzelman):

The lines A and B are prependicular to each other. And you know that the slope of line A is -1/5. CAn you then tell me: What is the slope of line B?

OpenStudy (mouth4war):

idk. i dont understand

OpenStudy (solomonzelman):

I will try to sa the same thing again, but in a more clear why

OpenStudy (solomonzelman):

Suppose you have two lines C and D, and suppose that you are given that lines C and D are perpendicular to each other. Now, say `the slope of line C is: M` Then, knowing that line D is prependicular to line C, `the slope of line D must be: -1/M `

OpenStudy (mouth4war):

the slope of line D is 5 but i dont understand how thats relevant to what i need to solve?

OpenStudy (solomonzelman):

You have two lines A and B that are perpendicular to each other, and thus you know that if the slope of line A is: M Then, the slope of line B is: -1/M +++++++ Now, if the slope of line A is -1/5 (that is our actual slope of line A) Then, the slope of line B is ?

OpenStudy (mouth4war):

0.2?

OpenStudy (solomonzelman):

\(\LARGE \color{black}{{\rm m_{_B}}=\frac{-1}{\color{black}{{\rm m_{_A}}}}}\) where: \(\LARGE \color{black}{{\rm m_{_B}}}\) denotes the slope of line B. \(\LARGE \color{black}{{\rm m_{_A}}}\) denotes the slope of line A.

OpenStudy (solomonzelman):

(This is provided line A & B are prependicular, which is in fact the case)

OpenStudy (solomonzelman):

\(\LARGE \color{black}{{\rm m_{_B}}=\frac{-1}{\color{black}{{\rm -0.2}}}=?}\)

OpenStudy (mouth4war):

i dont understand. thanks for trying to help me though but i dont think ill ever understand it

OpenStudy (solomonzelman):

Sorry that I can't help much... yw tho